# Effects of galactic environment on size and dark matter content in low-mass galaxies

### v1

@@ 1 Introduction
Other works focus on studying how galactic environment affects the SMR (Cooper et al., 2012; Cappellari, 2013; Lani et al., 2013; Cebrián & Trujillo, 2014; Delaye et al., 2014; Shankar et al., 2014; Kelkar et al., 2015; Zhang & Yang, 2019; Yoon et al., 2017, 2023). Specifically, Yoon et al. (2017) show that the most massive early-type galaxies tend to be larger in dense environments, likely as a consequence of the numerous mergers they experience throughout their lifetimes. This is in line with the idea that the size of massive dispersion-supported systems may be inflated over time if mergers impart sufficient energy (Boylan-Kolchin et al., 2006; Kaufmann et al., 2007; Trujillo et al., 2011).
@@
Another property closely linked to galaxy formation and evolution is the content of dark matter within the galaxy (Dutton et al., 2010; Auger et al., 2010; Kravtsov, 2024; Benavides et al., 2024). In particular, Kravtsov (2024) shows that Milky Way satellites and some UDGs follow a relation between the dark matter content within a galaxy’s half mass radius ($M^{50}_{\rm DM}$) and its stellar mass. Understanding how the SMR and the $M^{50}_{\rm DM}$ $M_{\star}$ relation depend on galactic environment for a broad range of low-mass galaxies is especially important as observatories like the Vera C. Rubin Observatory and Nancy Grace Roman Space Telescope usher in a new era of astronomy and give us access to a statistical samples low-mass galaxies across all environments.
@@ 2.2 Halo finding
We identify and characterize dark matter halo properties with the AMIGA Halo Finder [^2] http://popia.ft.uam.es/AHF/ (AHF; Knollmann & Knebe, 2009). We define dark matter halos to be spherical systems with viral radii, $r_{\rm vir}$, inside which the average density is equal to $\Delta(z)\rho_{\rm m}(z)$. The virial mass is then defined as the mass within $r_{\rm vir}$,
@@
For satellites, we take $r_{\rm halo}$ to be the AHF-defined truncated radius ($r_{\rm t}$), reflecting the subhalo’s reduced extent due to tidal stripping and environmental effects, and $M_{\rm halo}$ to be the mass within this truncated radius:
$r^{\rm satellites}_{\rm halo}\equiv r_{\rm t},\,\,{\rm and}\,\,M^{\rm satellites}_{\rm halo}\equiv M_{\rm DM}(<\,r_{\rm t}).$ (3)
@@ 2.4 Quantifying environment
For the purposes of this paper, we adopt the Perturbation Index (PI) to quantify galaxy environment (Dahari, 1984; Verley et al., 2007; Choi et al., 2018; Martin et al., 2019; Jackson et al., 2021). PI offers several advantages over other environmental indicators. Specifically, it not only directly measures the tidal influence of nearby companions by accounting for both their mass and proximity but also incorporates the properties of the primary galaxy, such as its mass and size. This provides a more comprehensive assessment of environmental impact. Additionally, PI’s ability to quantify the strength of tidal interactions provides a direct link to the physical processes (e.g., tidal stripping and mergers) that can significantly alter a galaxy’s size and dark matter distribution.
@@
where $M^{p}_{\rm halo}$ and $r^{p}_{\rm halo}$ are the halo mass and radius of the primary galaxy, $M^{c}_{\rm halo}$ is the halo mass of a companion galaxy $c$, and $D_{cp}$ is the 3D distance from the primary to its companion galaxy. This formulation of the PI estimates the cumulative tidal field about a galaxy due to the nearest $k$ neighbors out to a maximum radius, $r_{\rm max}$. We calculate the PI for each galaxy in our sample, taking into account the total contribution of all galaxies out to a maximum radius of $r_{\rm max}=\sqrt{3}L_{\rm box}/2$ in order to account for the effects from all galaxies within simulation box. We do this while appropriately dealing with the simulation box’s periodic boundary conditions.
Setting $r_{\rm max}$ to smaller values (e.g. 8 Mpc or 1 Mpc) only results in slightly lower values of PI, and this only affects the most isolated galaxies. Additionally, we note that we use halo masses and radii in this formulation, rather than the stellar mass and galactic size ($r_{50}$), given that tidal effects are certainly important out to $r_{\rm halo}$. This results in a distribution of PI’s that are more closely centered around unity, such that a PI $\gg$ 1 represents a galaxy that is highly perturbed out to its halo radius by its surroundings (i.e. $F_{\rm tidal}>F_{\rm bind}$).
@@ 2.5 Random forest analysis
Given the inherently complex nature of galaxy formation and evolution, it is plausible that a number of processes influence a galaxy’s size and inner dark matter content. To better capture the intricacy of these relationships, we adopt a machine learning approach that allows us to explore our data without confining the analysis to a search for simple power-laws. Specifically, we utilize a Random Forest (RF) regressor (Breiman, 2001). Unlike methods such as Principal Component Analysis (PCA; Pearson, 1901; Jolliffe, 2002) or autoencoders (Rumelhart et al., 1988), which are primarily used for dimensionality reduction or feature extraction, a Random Forest can directly model nonlinear relationships between features and the target variables, and crucially, provide a measure of feature importance. This makes it well-suited for our goal of identifying which global galactic and environmental properties most significantly determine a galaxy’s relative extent and relative inner dark matter content.
In this analysis, we focus on three key features that are physically relevant to galaxy formation and structure: the halo mass ($M_{\rm halo}$), the stellar-to-halo mass ratio ($M_{\star}$/$M_{\rm halo}$), and the Perturbation Index (PI). Halo mass sets the galaxy’s gravitational framework and regulates the distribution of baryons and dark matter. The stellar-to-halo mass ratio provides insight into star formation efficiency over a galaxy’s history. The Perturbation Index encodes environmental factors, such as tidal forces and interactions, that can reshape a galaxy’s structure. By training our model on $M_{\rm halo}$, $M_{\star}$/$M_{\rm halo}$, and PI, we aim to isolate the most influential drivers behind deviations in galaxy sizes and inner dark matter content from median scaling relations.
@@ 2.5.1 Overview
Random Forest regression is an ensemble learning method that constructs a multitude of decision trees during training and outputs the average prediction of the individual trees to improve predictive accuracy and reduce variance (Breiman, 2001). Each tree is built on a bootstrap sample, which is created by drawing random samples from the training set with replacement. This results in each tree being trained on a slightly different subset of the data. This bootstrap aggregation (or “bagging”) ensures that the trees are not identical, increasing the model’s robustness and reducing over-fitting. At each split in a tree, a subset of features is randomly selected for splitting, which introduces further diversity among trees. By averaging the predictions of these many randomized trees, the Random Forest produces a stable prediction that is less sensitive to the peculiarities of a single training subset.
While the primary function of a Random Forest regressor is to model the relationship between input features and target variables, our main objective in this study is to utilize it as a tool to assess the relative importance of various galactic and environmental features in predicting deviations from median scaling relations. Given our relatively small sample size (N $\sim$ 1200), creating a model that performs exceptionally well in terms of predictive accuracy is challenging. The limited dataset may not capture the full complexity of the underlying astrophysical processes, potentially hindering the model’s ability to generalize. Therefore, we focus on feature importance rather than maximizing predictive performance metrics such as $R^{2}$, defined as
@@
where $y_{i}$ are the true target values, $\hat{y}_{i}$ are the model predictions, and $\bar{y}$ is the mean of the true target values. A higher $R^{2}$ indicates that more variance in the target is explained by the model, but for the purposes of this analysis, it serves more as a reference than a goal. By emphasizing the identification of influential features, we aim to gain insights into the underlying physical processes that influence galaxy properties, despite the constraints imposed by the sample size.
@@ 2.5.2 Hyperparameters, data preparation and model training
The RF algorithm has several important hyperparameters that control the complexity and diversity of the ensemble:
  • n_estimators: The number of trees in the forest. A larger number typically yields more stable predictions but increases computational cost.
@@
  • max_features: The number (or fraction) of features to consider at each split. Controlling this affects the diversity of trees in the ensemble.
  • min_samples_split: The minimum number of samples required to split an internal node. Higher values make trees more conservative in splitting.
To mitigate the effects of outliers and to ensure that every feature contributes equally to the analysis, we preprocess the feature data using the RobustScaler from the scikit-learn.preprocessing module (Pedregosa et al., 2011). This scaler removes the median and scales the data according to the interquartile range (IQR), making it robust to outliers. This scaling centers the features around their median and scales them based on the 25th and 75th percentiles, making the scaling robust to extreme values.
To identify suitable hyperparameters, we use GridSearchCV with five-fold cross-validation. Rather than optimizing the entire hyperparameter space, we focus solely on optimizing max_features and min_samples_leaf to maximize $R^{2}$. We fix n_estimators at 600 to ensure stable predictions without excessive computational cost and leave all other parameters at their default values to maintain simplicity. By concentrating on optimizing max_features and min_samples_leaf, we directly control tree complexity directly, tuning only the parameters most likely to influence our model’s bias-variance trade-off given our small dataset. This approach reduces the risk of over-fitting to our limited sample while still enabling some meaningful optimization of model complexity.
After selecting the optimal hyperparameters, we apply them in all subsequent analyses. To ensure the robustness of our findings and that results are not dependent on any single train-test split, we repeat the random train-test splitting process 500 times, each time training the Random Forest regressor on 50% of the data and validating on the other 50%. This procedure yields distributions of performance metrics and feature importance scores, allowing us to assess their stability and variability.
@@ 2.5.3 Feature importance analysis
Feature importance in Random Forest regression provides a quantitative measure of the contribution of each feature to the model’s predictions. This is typically calculated by evaluating how much splitting on each feature reduces impurity on average. In our analysis, after training the model in each of the 500 iterations, we extract the feature importance scores, producing distributions of importance scores for each feature. This method offers a comprehensive understanding of which properties – halo mass, stellar-to-halo mass ratio, or environment – dominate in shaping deviations in galaxy sizes and inner dark matter content from the median scaling relations.
@@ 3.3 Random forest regression
After training the Random Forest regressor using the optimal hyperparameters determined via GridSearchCV (as listed in Table 1), we evaluate the model’s performance on the validation sets across the 500 iterations. For the relative extent ($\Delta\,\log\,r_{50}$), the mean $R^{2}$ score on the validation sets was approximately 0.32, indicating that the model explains approximately 32% of the variance in the deviations of the half-mass radius from the median size–mass relation for centrals at fixed stellar mass. For the relative inner dark matter content ($\Delta\,\log\,M^{50}_{\rm DM}$), the mean $R^{2}$ score was approximately 0.41, suggesting that the model accounts for about 41% of the variance in the deviations of the dark matter mass within $r_{50}$from the median $M^{50}_{\rm DM}$ – $M_{\star}$ relation for centrals at fixed stellar mass. While these $R^{2}$ values indicate moderate predictive power, our primary focus is on the relative importance of the features rather than the overall predictive accuracy.
Target max_features min_samples_leaf $\langle R^{2}\rangle$
Variable
$\Delta\,\log\,r_{50}$ ‘sqrt’ 3 0.32
$\Delta\,\log\,M^{50}_{\rm DM}$ ‘sqrt’ 3 0.41
Table 1: Optimal hyperparameters and performance. The columns show the optimized hyperparameters and average $R^{2}$ values for the two target variables ($\Delta\,\log\,r_{50}$ and $\Delta\,\log\,M^{50}_{\rm DM}$) over the 500 iterations.
Figure 6 shows the relative parameter importance in predicting relative extent ($\Delta\,\log\,r_{50}$; blue bars) and relative inner dark matter content ($\Delta\,\log\,M^{50}_{\rm DM}$; green bars), extracted from our Random Forest regressor model. The bar heights represent the median importance for each parameter, while the error bars mark the 5th and 95th percentiles of the importance distribution created by repeating the experiment 500 times. The input features are listed on the horizontal axis.

### v2

@@ 1 Introduction
Other works focus on studying how galactic environment affects the SMR (Cooper et al., 2012; Cappellari, 2013; Lani et al., 2013; Cebrián & Trujillo, 2014; Delaye et al., 2014; Shankar et al., 2014; Kelkar et al., 2015; Zhang & Yang, 2019; Yoon et al., 2017, 2023; Ghosh et al., 2024). Specifically, Yoon et al. (2017) show that the most massive early-type galaxies tend to be larger in dense environments, likely as a consequence of the numerous mergers they experience throughout their lifetimes. More recently, Ghosh et al. (2024) used a large sample of galaxies from the Hyper Suprime-Cam (HSC) survey to show a similar size-environment correlation at intermediate stellar masses, further reinforcing the idea that environmental influences extend across different mass regimes. This is in line with the idea that the size of massive dispersion-supported systems may be inflated over time if mergers impart sufficient energy (Boylan-Kolchin et al., 2006; Kaufmann et al., 2007; Trujillo et al., 2011).
@@
Another property closely linked to galaxy formation and evolution is the content of dark matter within the galaxy (Dutton et al., 2010; Auger et al., 2010; Romeo et al., 2020; Kravtsov, 2024; Benavides et al., 2024). The stellar-to-halo mass ratio ($M_{\star}$/$M_{\rm halo}$) provides key insight into the efficiency of star formation across different galaxy populations, with several studies exploring how it evolves with morphology, mass, and environment (e.g., Romeo et al., 2020). In particular, Kravtsov (2024) shows that Milky Way satellites and some UDGs follow a relation between the dark matter content within a galaxy’s half-mass radius ($M^{50}_{\rm DM}$) and its stellar mass. However, large-scale simulations often struggle to reproduce these relations consistently, highlighting challenges in accurately modeling the galaxy-halo connection (Romeo et al., 2020). Understanding how the SMR and the $M^{50}_{\rm DM}$ $M_{\star}$ relation depend on galactic environment for a broad range of low-mass galaxies is especially important as observatories like the Vera C. Rubin Observatory and Nancy Grace Roman Space Telescope usher in a new era of astronomy and give us access to statistical samples of low-mass galaxies across all environments.
@@ 2.1 The FIREbox Simulation Suite
While FIREbox has a modest cosmological volume compared to some larger-scale hydrodynamical simulations (Schaye et al., 2015; Crain et al., 2015; Pillepich et al., 2018; Nelson et al., 2018; Springel et al., 2018), it still spans a wide range of galactic environments, from low-density voids to galaxy groups. The simulation contains halos with virial masses up to $\sim 10^{13}$ $M_{\odot}$, capturing environments comparable to small galaxy groups, though it does not fully sample the most massive cluster environments. Nevertheless, FIREbox offers an important advantage over larger-volume simulations: its ability to resolve the interstellar medium (ISM) at high spatial resolution. This allows for a more realistic treatment of star formation and feedback processes, which are crucial for understanding the structural and dynamical properties of low-mass galaxies. Given our focus on low-mass galaxies and the importance of resolving their internal baryonic processes, FIREbox is well-suited for investigating how environment influences galaxy sizes and dark matter content.
@@ 2.2 Halo finding
We identify and characterize dark matter halo properties with the AMIGA Halo Finder [^2] http://popia.ft.uam.es/AHF/ (AHF; Knollmann & Knebe, 2009). We define dark matter halos to be spherical systems with viral radii, $r_{\rm vir}$, inside which the average density is equal to $\Delta(z)\rho_{\rm m}(z)$. The virial mass is then defined as the total mass within $r_{\rm vir}$,
@@
For satellites, we take $r_{\rm halo}$ to be the AHF-defined truncated radius ($r_{\rm t}$), reflecting the subhalo’s reduced extent due to tidal stripping and environmental effects, and $M_{\rm halo}$ to be the total mass ($M_{\rm tot}=M_{\rm DM}+M_{\rm bary}$) within this truncated radius:
$r^{\rm satellites}_{\rm halo}\equiv r_{\rm t},\,\,{\rm and}\,\,M^{\rm satellites}_{\rm halo}\equiv M_{\rm tot}(<\,r_{\rm t}).$ (3)
@@ 2.4 Quantifying environment
For the purposes of this paper, we adopt the Perturbation Index (PI) to quantify galaxy environment (Dahari, 1984; Verley et al., 2007; Choi et al., 2018; Martin et al., 2019; Jackson et al., 2021). The PI offers several advantages over other environmental indicators. Specifically, it not only directly measures the tidal influence of nearby companions by accounting for both their mass and proximity but also incorporates the properties of the primary galaxy, such as its mass and size. This provides a more comprehensive assessment of environmental impact. Additionally, the PI’s ability to quantify the strength of tidal interactions provides a direct link to the physical processes (e.g., tidal stripping and mergers) that can significantly alter a galaxy’s size and dark matter distribution.
@@
where $M^{p}_{\rm halo}$ and $r^{p}_{\rm halo}$ are the halo mass and radius of the primary galaxy, $M^{c}_{\rm halo}$ is the halo mass of a companion galaxy $c$, and $D_{cp}$ is the 3D distance from the primary to its companion galaxy. This formulation of the PI estimates the cumulative tidal field about a galaxy due to the nearest $k$ neighbors out to a maximum radius, $r_{\rm max}$. We calculate PI for each galaxy in our sample, summing contributions from all galaxies out to a maximum radius of $r_{\rm max}=\sqrt{3}L_{\rm box}/2$ to account for interactions throughout the full simulation volume while handling periodic boundary conditions.
We use halo masses and radii in our formulation, rather than the stellar mass and galactic sizes ($r_{50}$), given that tidal effects are certainly important out to $r_{\rm halo}$. This choice results in a PI distribution that is more naturally centered around unity, such that a PI $\gg$ 1 represents a galaxy that is highly perturbed by its surroundings (i.e. $F_{\rm tidal}>F_{\rm bind}$).
To assess the physical scale that PI probes, we compare PI values calculated using smaller values of $r_{\rm max}$ and find that for most galaxies, PI remains unchanged when restricting to $r_{\rm max}$ = 10 Mpc , but decreases when limiting to $r_{\rm max}$ = 1 Mpc. This suggests that PI is primarily sensitive to tidal interactions on scales of 1–10 Mpc rather than sub-Mpc interactions. Consequently, while PI effectively captures local tidal effects, it does not directly probe large-scale cosmic structures, which have been shown to influence galaxy properties in other studies (e.g., Muldrew et al., 2012).
@@ 2.5 Random forest analysis
Given the inherently complex nature of galaxy formation and evolution, it is plausible that a number of processes influence a galaxy’s size and inner dark matter content. To better capture the intricacy of these relationships, we adopt a machine learning approach that allows us to explore our data without confining the analysis to a search for simple power-laws. Specifically, we utilize a Random Forest (RF) regressor (Ho, 1995).
An RF regressor can directly model nonlinear relationships between input parameters (commonly referred to as “features”) and the output quantities of interest (known as “target variables”). Features are the independent variables used as inputs to the model, while target variables represent the dependent quantities the model is trained to predict. In our case, the features correspond to global galactic and environmental properties, while the target variables quantify deviations in galaxy sizes and inner dark matter content from their expected median relations.
An important advantage of this approach is that Random Forest regression not only predicts these outcomes but also provides a measure of feature importance, allowing us to identify which properties most significantly influence a galaxy’s structural and dark matter characteristics.
In this analysis, we focus on three features that are physically relevant to galaxy formation and structure: the halo mass ($M_{\rm halo}$), the stellar-to-halo mass ratio ($M_{\star}$/$M_{\rm halo}$), and the Perturbation Index (PI). Halo mass sets the galaxy’s gravitational framework and regulates the distribution of baryons and dark matter on a cosmological scale. The stellar-to-halo mass ratio provides insight into star formation efficiency over a galaxy’s history. The Perturbation Index encodes environmental factors, such as tidal forces and interactions, that can reshape a galaxy’s structure. By training our model on $\log$($M_{\rm halo}$), $\log$($M_{\star}$/$M_{\rm halo}$), and $\log$(PI), we aim to isolate the most influential drivers behind deviations in galaxy sizes and inner dark matter content from median scaling relations.
@@ 2.5.1 Overview
Random Forest regression is an ensemble learning method that constructs a multitude of decision trees during training and outputs the average prediction of the individual trees to improve predictive accuracy and reduce variance (Breiman, 2001).
Each tree consists of a series of binary decision splits, where at each node, the algorithm selects a feature and a corresponding threshold (splitting criterion) that best separates the data. In our implementation, we use mean squared error (MSE) as the splitting criterion to determine the optimal split at each node. MSE measures the average squared difference between the true and predicted values, ensuring that larger errors contribute more significantly to the loss function.
Formally, MSE is computed as:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2},$ (6)
where $y_{i}$ are the target values, $\hat{y}_{i}$ are the model predictions, and N is the total number of samples.
The splitting process continues recursively until a stopping condition is met, such as reaching a predefined maximum tree depth or a minimum number of samples per leaf node. A leaf node is the final, unsplit node of the tree, which contains the predicted output value.
Each decision tree is trained on a bootstrap sample, meaning it is built using a randomly drawn subset of the training data (with replacement). This bootstrap aggregation (or “bagging”) ensures that each tree is trained on a slightly different dataset, increasing robustness and reducing sensitivity to individual data points. Additionally, at each split, the algorithm only considers a randomly selected subset of available features, further promoting variation among the trees and improving model generalization.
While the primary function of a Random Forest regressor is to model the relationship between input features and target variables, our main objective in this study is to utilize it as a tool to assess the relative importance of various galactic and environmental features in predicting deviations from median scaling relations. Given our relatively small sample size (N $\sim$ 1200), achieving high predictive accuracy is inherently challenging. The limited dataset may not fully capture the statistical diversity of the physical processes at play, potentially restricting the model’s ability to generalize.
Thus, instead of prioritizing predictive performance, we focus on feature importance. To quantify this, we use $R^{2}$ (coefficient of determination), defined as:
@@
where $\bar{y}$ is the mean of the true target values. A higher $R^{2}$ indicates that more variance in the target variable is explained by the model, but in this analysis, it primarily serves as a reference rather than an optimization goal. By emphasizing the identification of influential features, we aim to gain insights into the underlying physical processes that influence galaxy properties, despite the constraints imposed by the sample size.
@@ 2.5.2 Hyperparameters, data preparation and model training
The RF algorithm has several important hyperparameters that control the complexity of the model:
  • n_estimators: The number of trees in the forest. A larger number typically yields more stable predictions but increases computational cost. This value also saturates, so there is always a minimum number of trees for a given size of the dataset that will lead to stable predictions
@@
  • max_features: The number (or fraction) of features to consider at each split. Controlling this affects the diversity of trees in the ensemble and the degree to which each tree is required to explore the full set of features, rather than the best performing one..
  • min_samples_split: The minimum number of samples required for an internal node to be split. Increasing this value reduces the likelihood of creating deep, highly specialized trees by requiring more data to justify a split, thereby promoting broader, more generalized decision boundaries.
Before training the model, we preprocess our features using RobustScaler from the scikit-learn (Pedregosa et al., 2011). This scaler removes the median and scales the data according to the interquartile range (IQR), making it robust to outliers. This scaling centers the features around their median and scales them based on the 25th and 75th percentiles, making the scaling robust to extreme values.
We identify the optimal hyperparameters using GridSearchCV with five-fold cross-validation. Rather than optimizing the entire hyperparameter space, we focus solely on optimizing max_features and min_samples_leaf to maximize $R^{2}$. We fix n_estimators at 600 to ensure stable predictions without excessive computational cost and leave all other parameters at their default values to maintain simplicity. By concentrating on optimizing max_features and min_samples_leaf, we directly control tree complexity directly, tuning only the parameters most likely to influence our model’s bias-variance trade-off given our small dataset. This approach reduces the risk of over-fitting to our limited sample while still enabling some meaningful optimization of model complexity.
Although we use MSE as the RF criterion, we use $R^{2}$ in GridSearchCV because it provides a more intuitive measure of model performance. The two metrics are mathematically related: minimizing MSE maximizes $R^{2}$, since the numerator of the second term in Equation 7 is simply the total MSE. Additionally, $R^{2}$ is unitless, allowing for easier comparison across different target variables.
After selecting the optimal hyperparameters, we apply them in all subsequent analyses. To ensure the robustness of our findings and that results are not dependent on any single train-test split, we repeat the random train-test splitting process 500 times, each time training the Random Forest regressor on 50% of the data and testing on the other 50%. This procedure yields distributions of performance metrics and feature importance scores, allowing us to assess their stability and variability.
@@ 2.5.3 Feature importance analysis
Feature importance in Random Forest regression provides a quantitative measure of the contribution of each feature to the model’s predictions.The importance of a feature is computed as the total weighted reduction in MSE across all nodes where the feature is used for splitting.
Mathematically, for a feature $x_{j}$ , its importance score is given by:
$I(x_{j})=\sum_{\text{splits on }x_{j}}\frac{N_{\text{node}}}{N_{\text{total}}}\Delta\text{MSE},$ (8)
where $N_{\text{node}}$ is the number of samples at the node where the split occurs, and $\Delta$ MSE is the decrease in mean squared error due to the split.
After training the model in each of the 500 iterations, we extract the feature importance scores, producing distributions of importance scores for each feature. This method offers a comprehensive understanding of which properties – halo mass, stellar-to-halo mass ratio, or environment – dominate in shaping deviations in galaxy sizes and inner dark matter content from the median scaling relations.
@@ 3.3 Random forest regression
After training the Random Forest regressor using the optimal hyperparameters determined via GridSearchCV (as listed in Table 1), we evaluate the model’s performance on the test sets across the 500 iterations. For the relative extent ($\Delta\,\log\,r_{50}$), the mean $R^{2}$ score on the test sets was approximately 0.32, indicating that the model explains approximately 32% of the variance in the deviations of the half-mass radius from the median size–mass relation for centrals at fixed stellar mass. The corresponding mean MSE is approximately 0.1, providing a direct measure of the average squared error in the model’s predictions.
For the relative inner dark matter content ($\Delta\,\log\,M^{50}_{\rm DM}$), the mean $R^{2}$ score was approximately 0.41, suggesting that the model accounts for about 41% of the variance in the deviations of the dark matter mass within $r_{50}$ from the median $M^{50}_{\rm DM}$ – $M_{\star}$ relation for centrals at fixed stellar mass. The mean MSE for $\Delta\,\log\,M^{50}_{\rm DM}$ is approximately 0.2, reflecting the typical magnitude of prediction errors. While these $R^{2}$ values indicate moderate predictive power, our primary focus is on the relative importance of the features rather than the overall predictive accuracy.
Target max min $\langle R^{2}\rangle$ $\langle{\rm MSE}\rangle$
Variable features samples leaf
$\Delta\,\log\,r_{50}$ ‘sqrt’ 3 0.32 0.10
$\Delta\,\log\,M^{50}_{\rm DM}$ ‘sqrt’ 3 0.41 0.20
Table 1: Optimal hyperparameters and performance. The table presents the optimized hyperparameters alongside the mean $R^{2}$ and MSE values for the two target variables ($\Delta\,\log\,r_{50}$ and $\Delta\,\log\,M^{50}_{\rm DM}$), averaged over 500 iterations.
Figure 6 shows the relative parameter importance in predicting relative extent ($\Delta\,\log\,r_{50}$; blue bars) and relative inner dark matter content ($\Delta\,\log\,M^{50}_{\rm DM}$; green bars), extracted from our Random Forest regressor model. [^4] Setting max_features=None instead of max_features=‘sqrt’ alters the absolute importance values but does not change the relative ranking of features. The bar heights represent the median importance for each parameter, while the error bars mark the 5th and 95th percentiles of the importance distribution created by repeating the experiment 500 times. The input features are listed on the horizontal axis.